OpenStudy (anonymous):

what is is the right equation for y+15=11

OpenStudy (anonymous):

what number would you subtract from 15 to get 11

OpenStudy (anonymous):

y = 11 - 15

OpenStudy (anonymous):

yeah thats correct

OpenStudy (anonymous):

wht its not

OpenStudy (anonymous):

you must subract 15 from both sides you get y + 0 = 11 - 15